how do you completely factor 10x^2-5x-5
Answer by ewatrrr(24785)

 
Hi,
10x^2-5x-5 = 5(2x^2 -x-1)= 5(2x+1)(x-1)
check factoring of (2x^2 -x-1) using FOIL
F First terms
O Outside terms
I Inside terms
L Last terms